Label HMNR peaks and indenify structure

(a) HN (2) CH N (6) CHAN

a) C8H11N
b) C8H11N
c) C7H9N
Answer #1

Questions :- 1-HNMRs :- @) Co M, N :-As per molecular formula, calculate DBE DBE( Double Bond Equivalence) = C+1+H+X-N WhereDBE by Molecular frarossa, DBE = 8+1- (6) COMIN &- Calculating c+ 2 H+X-N) 4-69 70-4 - 9-5 IDBE = 4 It meary that there are 4